OpenStudy (anonymous):

a sail is in the form of a right triangle that is 3 times as high as it is wide. the sail is made from 6 square meters of material. what is the height

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

we need a variable for the width. what do you want to call it?

OpenStudy (anonymous):

Area of a Triangle = 1/2 * bh The base can be x and the height 3x So, 6 = 1/2 * x(3x) 6 = 1/2 * 3x^2 12 = 3x^2 4 = x^2 x = 2

OpenStudy (anonymous):

i usually use x

OpenStudy (misty1212):

ok then, but @sunnyshores pretty much gave you the complete answer using \(x\) as the width, so your answer is above

OpenStudy (anonymous):

thanks for helping